Circular No. 7707

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1650–500 R. Remillard, reporting for the RXTE All-Sky Monitor (ASM) team at

the Massachusetts Institute of Technology and Goddard Space Flight Center (GSFC), NASA, writes: “The RXTE ASM has detected x-ray emission from a new source, XTE J1650–500. The ASM position is centered at α = 16^{h}50^{m}00^{s}, δ = −50^{o}00′ (equinox 2000.0; estimated uncertainty 6′). The first detection was made on Sept. 5.637 UT, with an x-ray intensity (2–12 keV) of 0.14(2) Crab, rising to 0.24(1) Crab on Sept. 6.432 and 0.27(1) Crab on Sept. 6.896. The ASM detections in the three energy bands suggest x-ray colors consistent with a ‘soft x-ray transient’, rather than an x-ray pulsar. Observations at radio, optical, and infrared wavelengths are needed as soon as possible.”

C. Markwardt, J. Swank, and E. Smith, GSFC, report on behalf of the RXTE team at GSFC: “The transient x-ray source XTE J1650–500 discovered with the RXTE ASM was confirmed with PCA measurements on Sept. 6.39 (0.33 Crab) and 6.62 UT (0.39 Crab). PCA scans in the second observation give a best fitting α = 16^{h}50^{m}00^{s}, δ = −49^{o}59′ (equinox 2000.0; estimated uncertainty ∼ 2′). The continuum spectrum was a power law with a photon index of 1.5, plus a blackbody with temperature kT = 1.2 keV (∼ 10 percent of total), which is typical of hard states of black-hole candidates. No pulsations were seen.”

A. J. Castro-Tirado, Instituto de Astrofísica de Andalucía, Granada; P. Kilmartin, A. Gilmore, O. Petterson, and I. Bond, Mt. John University Observatory; P. Yock, Auckland University; and C. Sánchez-Fernández, Laboratorio de Astrofísica Espacial y Física Fundamental, Madrid, communicate: “We have observed the RXTE error box of the probable soft x-ray transient XTE J1650–500 with the Mt. John Observatory 0.6-m telescope at Lake Tekapo, New Zealand. A 180-s image obtained with a wide blue-band filter centered on the RXTE position reveals a possible candidate of mag ∼ 17 (preliminary) that might be present as a faint object (near the detection limit) in the corresponding blue Palomar Sky Survey image. A preliminary position is α = 16^{h}50^{m}01^{s}.0, δ = −49^{o}57′45′′ (equinox 2000.0; ± 2′′), which is nearly 1′ south of the bright star HDE 151476 (V = 8.0). Spectroscopic observations are encouraged in order to confirm whether this is the optical counterpart to XTE J1650–500. An identification chart is

posted at http://www.iaa.es/~ajct/XRBs/xtej1650.html.”

2001 September 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green